# pacer: Preference-conditioned All-terrain Costmap Generation

### v1

@@ VI-A Aerial Map Experiments
In our simulated experiments, we build aerial maps from drone footage of three locations around our campus, which we consider seen environments. We also use open-source aerial maps [23] from around the world, covering a wide variety of both urban and natural terrain types and which we consider unseen. For each of the seen and unseen environments, we provide a start and goal location and test varying operator preferences. We test realistic preferences, and “inverted” preferences, in which each of the pairwise orderings in the realistic preference are reversed. Given the robot’s pose on the aerial map, the robot’s projected bird’s eye view can be found, and used as input to generate the local costmap. Planning is done using the A* algorithm [24] on the costmap.
Table II displays the results for seen environments. Towards answering question 1a, pacer has similar results to the sterling baseline, as both approaches were trained on the same in-distribution data. When no useful context has been provided for pacer (i.e. the context contains only unseen terrains which do not match the environment as per 1b), the results are on par with the classifier baseline. pacer’s success despite the lack of an informative preference context shows that the model has captured a prior over realistic cost-assignment for in-distribution terrains.
Table II displays the results for unseen environments. Towards answering question 2a, when given an informative context, pacer outperforms the sterling baseline. Though representation-learning approaches like sterling should theoretically generalize due to their continuous representation space, this is contingent on similar terrains forming clusters in this space, which may not be the case for unseen terrains. Note that here, for unseen environments, the classifier baseline has been omitted, as it allows no way for a user to provide terrain preferences for classes that are not predefined. pacer overcomes the limitations of previous paradigms, as it both allows preferences in unseen environments to be expressed and generalizes well to these unseen environments. Additionally, when no useful context has been provided for the unseen terrains (per 2b), pacer is not able to adapt.
@@
Low (%) Medium (%) High (%)
@@
Uninformative Context
$\mathcal{D}_{\mathrm{real}}$ 66.22% 14.48% 19.30%
$\mathcal{D}_{\mathrm{shuffled}}$ 81.82% 7.67% 10.51%
$\mathcal{D}_{\mathrm{synthetic}}$ 67.97% 15.89% 16.15%
TABLE III: Proportion of planned paths that traverse low, medium, and high-cost terrains in seen environments for each training phase. The model after each phase of training is tested using realistic, inverted, and uninformative contexts
Method Low (%) Medium (%) High (%)
@@
No Context
$\mathcal{D}_{\mathrm{real}}$ 73.44% 7.50% 19.06%
$\mathcal{D}_{\mathrm{shuffled}}$ 82.35% 4.65% 13.00%
$\mathcal{D}_{\mathrm{synthetic}}$ 53.83% 2.20% 43.97%
TABLE IV: Proportion of planned paths that traverse low, medium, and high-cost terrains in unseen environments for each training phase. The model after each phase of training is tested using realistic, inverted, and uninformative contexts
@@ VI-B Ablation Study
To understand the effects of each phase in the training process, we perform an ablation study using the same environments as in the aerial simulator experiments.
In the $\mathcal{D}_{\mathrm{real}}$ phase, the model is trained only on real data and realistic preferences. In the $\mathcal{D}_{\mathrm{shuffled}}$ phase, the model is pretrained on real data with realistic preferences, and then trained on a smaller amount of changed preferences. In $\mathcal{D}_{\mathrm{synthetic}}$ phase, the model is trained according to all three phases.
Results are shown in Tab. IV, IV. In seen environments, the model trained on $\mathcal{D}_{\mathrm{real}}$ performs the best of the three with realistic preferences, but is unable to adapt when the preferences are inverted. Even without a useful context provided, the majority of planned trajectories are on low-cost terrain. In unseen environments, this method performs the worst regardless of preference context. The model trained on $\mathcal{D}_{\mathrm{shuffled}}$ is able to adapt to changing preference orderings in seen environments, though is unable to recognize and match new terrains in the preference context to the new environment. The model trained on $\mathcal{D}_{\mathrm{synthetic}}$ is shown to both respect changing preference order and recognize new terrains.
VI-C Discussion
@@ VII Real Robot Experiments
TABLE V: Number of successes per 5 trials of different approaches across various environments. A trial is a success if the robot reaches the goal without traversing across undesirable terrain and without operator intervention.
Table V shows results from real robot experiments. Our approach had the most successful trials across all environments. While the classifier performed well in environments 1 and 2, we hypothesize that difficult lightning conditions and variations in terrain appearance caused the failures in environments 3 and 4. Though sterling performed as the best baseline in the simulated experiments (which involved only planning), it seemed to be unable to execute these plans in real-robot experiments. Many of the failure cases involved the robot driving slightly off-path and just grazing the undesirable terrains. In patch-based representations, a single patch may contain multiple different terrains (e.g. half sidewalk and half grass), so the cost assigned to the patch would be some combination of the different terrain costs, resulting in a coarser degree of control on the physical robot. Our approach overcomes this limitation since it directly outputs a fine-grained costmap in a single forward pass.
@@ VIII Limitations and Future Work
The experiments reported in this paper were conducted on a single campus with one robot. Future work should extend the study to more varied terrains and deploy multiple robots with diverse sensors. Additionally, we aim to evaluate the ease with which human users can express preferences using our method, as well as explore alternative mechanisms for preference expression in costmap generation. Incorporating depth data into the costmap generation is another key area for improvement, as our current reliance on a homography transformation assumes flat ground, leading to limitations like the inability to avoid obstacles such as concrete curbs, despite their preferable terrain.
@@ X Acknowledgements
Appendix A Dataset Details
[…]
To construct a large training dataset from a relatively small amount of recorded data, we employ synthetic augmentation using 14 additional terrain textures, including sand, asphalt, leaves, river pebbles, cracked mud, and snow. Examples of synthetically augmented data are shown in Fig 6. Model training hyperparameters and statistics for the patch bank are shown in Tab. VII and Tab. VII.
Appendix B Aerial Map Simulator
In Fig 7 and 8, we show examples of paths planned using pacer in our aerial map simulator. These experiments are an evaluation purely over planning based on terrain preference, with no kinodynamic constraints, no errors due to localization, and no costs associated with elevation.
Given a robot state $(x,y,\theta)$ relative to the map, the robot’s view from this pose is found by overlaying the warped field-of-view shape onto the aerial map at that state. We employ A* on an eight-connected grid for path planning on the local costmap. If the node-expansion operation in A* explores a neighbor node which does not have an associated cost (meaning its state has not been captured in any previous view), the BEV from the current node is used to construct a local costmap which includes the previously-unobserved neighbor state, and the planning continues. Thus, the robot does not actually move its position in this path planning simulation. This is an optimization such that we can avoid transforming the full aerial map (which may be very large) into a costmap. Fig 7 shows the progression of the local costmaps as the robot discovers more of the environment and the resulting planned path.
Figure 7: Example of a costmap, planned path, and preference context of a simulated robot deployment in one of the natural unseen environments of the Aerial Map Experiments using our method. The costmap displays the timeseries progression of the local BEV costmaps generated with our method.
Appendix C Real Robot Experiments Setup
We run our real-robot experiments on a Clearpath Jackal Unmanned Ground Vehicle. A Microsoft Azure Kinect RGB-D camera supplies visual information at $\sim\!30\text{Hz}$. We run all algorithms on a Nvidia GeForce GTX 1050Ti GPU. Taking advantage of batched cost computation on the GPU, we find that the sterling and classifier baselines publish motion commands at $\sim\!20\text{Hz}$. Our approach publishes motion commands at $\sim\!11\text{Hz}$ on the same GPU since our approach outputs full costmaps, which is done sequentially.
Though each method was trained using data collected from a Boston Dynamics Spot, they were all deployed zero-shot on the Clearpath Jackal with only visual input. While most experimental results reported in sterling were obtained using a joint representation space with both visual and inertial-propriceptive-tactile features deployed on the legged Spot, the approach claims to be broadly adaptable to robots of different morphology with vision-only navigation. We follow the precedent set in sterling of transferring visual features learned from the Spot to the Jackal.
Figure 9: Example of homography-transformed BEV image of live robot deployment and corresponding local costmap. The best local plan is chosen from a set of path options based on a weighted combination of the terrain cost and distance to goal.
C-A Sampling-based Preference-aligned Local Planning Formulation
We formalize the local path planning problem as a search for the optimal motion arc $\Gamma^{*}$ from a set of arc options $\{\Gamma_{1},\Gamma_{2},\ldots,\Gamma_{n}\}$, each of which adheres to the kinodynamic constraints of Ackermann Steering (Fig 9). Each arc $\Gamma_{i}$ represents a feasible trajectory that the robot can follow, and is discretized into states/actions at finite future timestamps in a receding horizon planner. Each arc can be expressed as $\Gamma_{i}=\{s_{0},s_{1},\ldots,s_{m}\}$, where $s_{0}$ is the current state and $s_{m}$ is the state at the end of the planning horizon. More precisely, given a cost function $\mathcal{J}$, the optimal arc $\Gamma^{*}$ is
$\Gamma^{*}=\arg\min_{\Gamma}\mathcal{J}(\Gamma,G)\;,$
where
$\mathcal{J}(\Gamma_{i},G)=\alpha\cdot\mathcal{J}_{\text{geometric}}(\Gamma_{i},G)+(1-\alpha)\cdot\mathcal{J}_{\text{terrain}}(\Gamma_{i})\;,$
and $\alpha$ is the parameter representing the tradeoff between the geometric and terrain costs in the arc selection. $\mathcal{J}_{\text{geometric}}(\Gamma_{i},G)$ measures the geometric distance between the goal $G$ and the closest state in $\Gamma_{i}$. $\mathcal{J}_{\text{terrain}}(\Gamma_{i},G)$ is the terrain-cost of $\Gamma_{i}\in\{\Gamma_{1},\Gamma_{2},\ldots,\Gamma_{n}\}$, formulated as follows:
$\mathcal{J}_{\text{terrain}}(\Gamma_{i})=\sum_{s_{j}\in\Gamma_{i}}\gamma^{j}\cdot C(s_{j})\;,$
where the term $\gamma^{j}$ represents a decay factor applied to the cost of each state $s_{j}$ along the arc $\Gamma_{i}$ such that the planner prioritizes short-term navigational feasibility. In our experiments, we use $\gamma=0.9$. The cost at state $s_{j}$, $C(s_{j})$, is calcuated either by indexing into the full BEV costmap (as with pacer) or via terrain patch to scalar cost (as with the sterling and classifier baselines).
Having determined $\Gamma^{*}$, a finite action control sequence is obtained by applying 1D time optimal control along $\Gamma^{*}$, resulting in a series of control inputs that drive the robot along the optimal path.
Appendix D Preference Context Embeddings
We include a t-SNE visualization of the embedding space of the context encoder in Fig. 10. We display only the realistic preference contexts (red) and inverted preference contexts (blue), as the differences in other variations of preference context are much more subtle. The separation between realistic preference contexts (red) and inverted preference contexts (blue) indicate the encoder has learned useful features on which condition the output costmap.
Figure 10: Visualization of t-SNE plot showing embedding vectors of realistic preference contexts (red) and inverted preference contexts (blue).

### v2

@@ VI-A Aerial Map Experiments
In our simulated experiments, we build aerial maps from drone footage of three locations around our campus, which we consider seen environments. We also use open-source aerial maps [24] from around the world, covering a wide variety of both urban and natural terrain types and which we consider unseen. For each of the seen and unseen environments, we provide a start and goal location and test varying operator preferences. We test realistic preferences, and “inverted” preferences, in which each of the pairwise orderings in the realistic preference are reversed. Given the robot’s pose on the aerial map, the robot’s projected bird’s eye view can be found, and used as input to generate the local costmap. Planning is done using the A* algorithm [25] on the costmap. In Fig. 5, we show examples of paths planned using pacer in our aerial map simulator. These experiments are an evaluation purely over planning based on terrain preference, with no kinodynamic constraints, no errors due to localization, and no costs associated with elevation.
Tab. II displays the results for seen environments. Towards answering question 1a, pacer has similar results to the sterling baseline, as both approaches were trained on the same in-distribution data. When no useful context has been provided for pacer (i.e. the context contains only unseen terrains which do not match the environment as per 1b), the results are on par with the classifier baseline. pacer’s success despite the lack of an informative preference context shows that the model has captured a prior over realistic cost-assignment for in-distribution terrains.
Tab. II displays the results for unseen environments. Towards answering question 2a, when given an informative context, pacer outperforms the sterling baseline. Though representation-learning approaches like sterling should theoretically generalize due to their continuous representation space, this is contingent on similar terrains forming clusters in this space, which may not be the case for unseen terrains. Fapt install r-base-coreor unseen environments, the classifier baseline has been omitted, as it allows no way for a user to provide terrain preferences for classes that are not predefined. pacer overcomes the limitations of previous paradigms, as it both allows preferences in unseen environments to be expressed and generalizes well to these unseen environments. Additionally, when no useful context has been provided for the unseen terrains (per 2b), pacer is not able to adapt.
@@
In-Dist. Scenario Low (%) Medium (%) High (%)
@@
TABLE III: Proportion of planned paths that traverse low, medium, and high-cost terrains in seen environments for each training phase. The model after each phase of training is tested using realistic and inverted contexts.
Out-of-Dist. Scenario Low (%) Medium (%) High (%)
@@
TABLE IV: Proportion of planned paths that traverse low, medium, and high-cost terrains in unseen environments for each training phase. The model after each phase of training is tested using realistic and inverted contexts.
@@ VI-B Ablation Study
To understand the effects of each phase in the training process, we perform an ablation study using the same environments as in the aerial simulator experiments. In the $\mathcal{D}_{\mathrm{real}}$ phase, the model is trained only on real data and realistic preferences. In the $\mathcal{D}_{\mathrm{shuffled}}$ phase, the model is pretrained on real data with realistic preferences, and then trained on a smaller amount of changed preferences. In $\mathcal{D}_{\mathrm{synthetic}}$ phase, the model is trained according to all three phases.
Results are shown in Tab. IV, IV. In seen environments, the model trained on $\mathcal{D}_{\mathrm{real}}$ performs the best of the three with realistic preferences, but is unable to adapt when the preferences are inverted. Results for realistic preferences (black) are significantly better than those for inverted preference (blue) for this method in Tab. IV. In unseen environments (Tab. IV), this method performs the worst regardless of preference context. The model trained on $\mathcal{D}_{\mathrm{shuffled}}$ is able to adapt to changing preference orderings in seen environments, seen by comparing results for realistic preferences (black) to inverted preferences (blue) in Tab. IV. However, it is unable to recognize and match new terrains in the preference context to the new environment as evidenced by the drop in performance from Tab. IV to Tab. IV. The model trained on $\mathcal{D}_{\mathrm{synthetic}}$ is shown to both respect changing preference order and recognize new terrains. Our findings indicate that training the model to respond to preferences reduces performance with realistic preferences in seen environments, as $D_{\mathrm{shuffled}}$ and $D_{\mathrm{synthetic}}$ have slightly poorer performance than $D_{\mathrm{real}}$ with realistic but far better performance with inverted preferences. Furthermore, performance across all three methods decreases from seen to unseen since the scenario is harder, but $D_{\mathrm{synthetic}}$ has the least drop in performance. Variable preferences and unseen environments introduce greater complexity, requiring models to adapt and generalize. We prioritize these harder, practical scenarios, accepting a trade-off with peak accuracy in simpler settings.
VI-C Performance on Test Set
There are infinite costmaps which may satisfy a preference over terrains (given as a partial order of terrains), since the scale of the numerical costs do not matter as long as the ordering is followed. Therefore, rather that direct comparison with a “ground-truth” costmap, we report the margin ranking error of predicted costmaps using holdout set from each dataset in Tab. V. We sample the 500 points across the costmap, then for each pair of points, calculate the error as 0 if the relative magnitude of the two costs is correct. Otherwise the error is the difference between the two costs. We report the mean error for each category.
For each dataset, the model which was trained on that dataset had the lowest error. Across all datasets, $\mathcal{D}_{\mathrm{synthetic}}$ had the lowest or second-lowest error, and had therefore the overall best performance.
Test Dataset $\mathcal{D}_{\mathrm{real}}$ $\mathcal{D}_{\mathrm{shuffled}}$ $\mathcal{D}_{\mathrm{synthetic}}$
Model
$\mathcal{D}_{\mathrm{real}}$ 0.017 0.099 0.072
$\mathcal{D}_{\mathrm{shuffled}}$ 0.086 0.019 0.064
$\mathcal{D}_{\mathrm{synthetic}}$ 0.053 0.069 0.013
TABLE V: Margin ranking error for each of the three models on the test set of each of the three datasets. Lower error is better.
VI-D Discussion
@@ VII Real Robot Experiments
Tab. VI shows results from real robot experiments. Our approach had the most successful trials across all environments. While the classifier performed well in environments 1 and 2, we hypothesize that difficult lightning conditions and variations in terrain appearance caused the failures in environments 3 and 4. Though sterling performed as the best baseline in the simulated experiments (which involved only planning), it seemed to be unable to execute these plans in real-robot experiments. Many of the failure cases involved the robot driving slightly off-path and just grazing the undesirable terrains. In patch-based representations, a single patch may contain multiple different terrains (e.g. half sidewalk and half grass), so the cost assigned to the patch would be some combination of the different terrain costs, resulting in a coarser degree of control on the physical robot. Our approach overcomes this limitation since it directly outputs a fine-grained costmap in a single forward pass.
@@
TABLE VI: Number of successes per 5 trials of different approaches across various environments. A trial is a success if the robot reaches the goal without traversing across undesirable terrain and without operator intervention.
@@ VIII Limitations and Future Work
The experiments reported in this paper were conducted on a single campus with one robot. Future work should extend the study to more varied terrains and deploy multiple robots with diverse sensors. Additionally, we aim to evaluate the ease with which human users can express preferences using our method, as well as explore alternative mechanisms for preference expression in costmap generation. Incorporating a variable-length preference context would also be an interesting direction for future work, enabling greater flexibility and adaptability in representing user preferences. Exploring other model architectures could further enhance the system’s robustness and performance in diverse scenarios. Finally, incorporating depth data into the costmap generation is another key area for improvement, as our current reliance on a homography transformation assumes flat ground, leading to limitations like the inability to avoid obstacles such as concrete curbs, despite their preferable terrain.